OpenStudy (anonymous):

A ball is thrown upward from the top of an 80ft building so that its height in feet above the ground after t seconds is h(t)=80+64t-16tˆ2. QUESTION: find the average velocity over the interval from t = 0 s to t = 2 s CONFUSING.......

OpenStudy (rajee_sam):

Rate of change of position is your velocity

OpenStudy (rajee_sam):

Position is your height given toyou. dh/dt will give you your average velocity

OpenStudy (rajee_sam):

I mean it will give you your velocity function. Now substitute t = 0 and you will get one value for velocity, Then put t = 2 you will get another value for velocity. take average of the two. There is your average velocity

OpenStudy (anonymous):

\[h(0) = 80m; h(2) = 144m\] Time is 2 seconds. So... position traveled is 144-80 = 64m time taken = 2 seconds. Velocity is in meters per second so... Velocity = 64/2 = 32m/s

OpenStudy (anonymous):

thank you , I tried, and it the same answer...thxx.

OpenStudy (rajee_sam):

which way did you try?

OpenStudy (anonymous):

let 0 and 2 into the equation. But I didnt do the substitution